list all possible whole numbers dimensions for rectangles
that have an area of 18 square feet. which has the greatest perimeter





Answer by rothauserc(4718)

The factors of 18 are,
1, 18, 2, 9, 6, 3
a) 1*18 = 18 square feet and Perimeter(P) = 2+36 = 38 feet
b) 2*9 = 18 square feet and P = 4+18 = 22 feet
c) 3*6 = 18 square feet and P = 6+12 = 18 feet
a) has the greatest perimeter.